SourMunchkin7806:

Two forces with magnitudes of 150 and 100 pounds act on an object at angles of 40° and 170°, respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

SourMunchkin7806:

@dude

SourMunchkin7806:

@Vocaloid

Valirium:

@Comerpickles

Comerpickles:

81.78 degrees

SourMunchkin7806:

I need to show how i get it

SourMunchkin7806:

Step by Step.

Comerpickles:

Okay. To solve, first split each force into x and y components. F1 = 150 F1_x = 150 cos(40) F1_y = 150 sin(40) F1_x = 114.91 F1_y = 96.42 Now the second force: F2 = 100 F2_x = 100 cos(170) F2_y = 100 sin(170) F2_x = -98.48 F2_y = 17.36 Notice that Force 2's x (horizontal) component is negative. This means that it points to the left. If you draw out a diagram you'll see what I mean. Now to solve for the resultant force: FR = resultant force FR_x = F1_x + F2_x FR_y = F1_y + F2_y FR_x = 16.43 FR_y = 113.78 What we have just solved for are the x and y components of the resulting force when the two forces are added together. Now to find the magnitude and direction of the resultant force, use the following formulas. Magnitude: use the distance formula FR = sqrt[ (FR_x)^2 + (FR_y)^2 ] FR = sqrt[ 13215.83 ] FR = 114.96 Direction: remember that tangent of an angle gives its slope. Since we have the slope and need the angle, we use arctangent. Slope = rise / run = vertical / horizontal = FR_y / FR_x = 6.93 Angle = arctangent(6.93) = 81.78 degrees

SourMunchkin7806:

Did you copy this or?

SourMunchkin7806:

So i can make revisions lol

Comerpickles:

The step by step version yes because mainly I would need a graph to do this, which I do not have at the time

SourMunchkin7806:

Ok thank you bro life freaking saver aha

Comerpickles:

Not really XD